Air is compressed in a diesel engine from an initial pressure of 13 psia and a temperature of 120 F to one-twelfth of its original volume. Calculate the final temperature assuming the compression is adiabatic


1
Expert's answer
2021-11-26T10:28:36-0500

"P_1 = 13 \; psia \\

T_1 = 120 \; °F = 322.039 \; K \\

V_2 = \frac{V_1}{12}"

Adiabatic compression

"\frac{P_2}{P_1} = (\frac{V_1}{V_2})^γ \\

\frac{T_2}{T_1} = (\frac{V_1}{V_2})^{γ-1} \\

P_2 = P_1 \times (12)^{1.4} \\

= 13 \times (12)^{1.4} \\

P_2 = 421.499 \; psia \\

T_2 = T_1 \times (\frac{V_1}{V_2})^{γ-1} \\

= 322.039 \times (12)^{0.4} \\

= 870.123 \; K \\

T_2 = 1106.551 \; °F ≈ 1010 \; °F"